OpenStudy (kenshin):

Can someone help me with htis epsilon delta proof please? My work in comments:

OpenStudy (kenshin):

1. The problem statement, all variables and given/known data Use the epsilon delta definition to show that lim(x,y) -> (0,0) (x*y^3)/(x^2 + 2y^2) = 0 2. Relevant equations sqrt(x^2) = |x| <= sqrt(x^2+y^2) ==> |x|/sqrt(x^2+y^2) <= 1 ==> |x|/(x^2+2y^2)? 3. The attempt at a solution This limit is true IFF for all values of epsilon > 0, there exists a delta such that: 0<sqrt(x^2+y^2)<delta ==> |(x*y^3)/(x^2+2y^2)| < epsilon When checking my answers, I'm stuck at this step: Can: |x|/sqrt(x^2+2y^2) <= 1 imply that |x|/(x^2+2y^2) will also be <= 1? I'm skeptical as it seems wrong, but if this step holds, then the rest of my proof can be as follows: (|x|/(x^2+2y^2))*|y^3| <= |y^3| = |(y^2)^(3/2)| <= (x^2 + y^2 ) ^ (3/2) < (delta)^3 thus let delta = (epsilon)^(1/3) and we get: sqrt(x^2+y^2)<epsilon^(1/3) ==> (x^2+y^2)^(3/2) < epsilon ==> (y^2)^(3/2) = y^3 < (x^2+y^2)^(3/2) < epsilon; Since |x|/(x^2+2y^2) <= 1, (|x|/(x^2+2y^2)) * |y^3| < |y^3| < epsilon; and we have f(x) < epsilon.

OpenStudy (michele_laino):

hint: if we rewrite such function, with polar coordinate, we get: \[\Large f\left( {\rho ,\theta } \right) = {\rho ^2}\cos \theta {\left( {\sin \theta } \right)^3}\]

OpenStudy (michele_laino):

now, we can apply the epsilon-delta method for \(\rho\) variable, being \(\theta\) an established value

OpenStudy (kenshin):

Ok thanks, I gotta go review my polar function chapters then.

OpenStudy (michele_laino):

sorry, I have misunderstood the function, please ignore my reply above

OpenStudy (michele_laino):

here is the right formula: \[\Large f\left( {\rho ,\theta } \right) = \frac{{{\rho ^2}\cos \theta {{\left( {\sin \theta } \right)}^3}}}{{1 + {{\left( {\sin \theta } \right)}^2}}}\]

OpenStudy (michele_laino):

now, for fixed \(\theta\), \(f\) is a function of \(\rho\) only

OpenStudy (michele_laino):

so, when \(x,y \to 0\), then \(\rho \to 0\), being \(\rho= \sqrt {(x^2+y^2)}\)

OpenStudy (kenshin):

Sorry, I haven't reviewed polar coordinates yet and can't follow this. but there's another hint I got from another forum: use (1/(x^2+2y^2)) <= 1/(2y^2) he said

OpenStudy (michele_laino):

yes! In that case we have to show that the new function \(g(x,y)\) checks this condition: \[\Large \mathop {\lim }\limits_{x,y \to 0} g\left( {x,y} \right) = 0\]

OpenStudy (michele_laino):

where: \[\Large g\left( {x,y} \right) = \frac{{x{y^3}}}{{2{y^2}}} = \frac{{xy}}{2}\]

OpenStudy (michele_laino):

now, using the polar coordinates, we have: \[\Large g\left( {\rho ,\theta } \right) = {\rho ^2}\frac{{\sin \left( {2\theta } \right)}}{4}\]

OpenStudy (michele_laino):

so, we can write this: \[\Large 0 \leqslant g\left( {\rho ,\theta } \right) \leqslant \frac{{{\rho ^2}}}{4}\] and for \(\rho \to 0\), we get \(g \to 0\)

OpenStudy (kenshin):

sweet, i'll review this question once i've done my reading, so it probably will take some time before I can understand and absorb all this, but thanks alot anyway!

OpenStudy (michele_laino):

:)

OpenStudy (michele_laino):

oops.. we can write this: \[ \Large - \frac{{{\rho ^2}}}{4} \leqslant g\left( {\rho ,\theta } \right) \leqslant \frac{{{\rho ^2}}}{4}\] and when \(\rho \to 0\) then \(g \to 0\)

OpenStudy (michele_laino):

since: \[\Large - 1 \leqslant \sin \left( {2\theta } \right) \leqslant 1\]

OpenStudy (kenshin):

does that mean I should set delta = 2sqrt(epsilon) here then? We need 0<p<delta to imply g(p,theta)< epsilon, since we know g(p,theta)< p^2/4, if we set delta = 2sqrt(epsilon), 0<p<2sqrt(epsilon) ==> (p^2)/4 < epsilon, then it works?

OpenStudy (michele_laino):

if we consider the function \(g\) then you are right, it is suffice to take \(\delta = \sqrt 2 \epsilon\), furthermore, we have to keep in mind that we are using the euclidean metric, so the neighbour about a point x is given by a circle centered at x

OpenStudy (kenshin):

but since $$1/(x^2+2y^2) <= 1/2y^2$$ therefore $$xy^3/(x^2+2y^2) <= g(p,theta)$$, would it be ok to then use the g function's delta for f?

OpenStudy (michele_laino):

I'm thinking...

OpenStudy (kenshin):

Ok I'll be right back after dinner thanks!

OpenStudy (michele_laino):

we can write this: \[\large \begin{gathered} f\left( {\rho ,\theta } \right) = \frac{{{\rho ^2}\cos \theta {{\left( {\sin \theta } \right)}^3}}}{{1 + {{\left( {\sin \theta } \right)}^2}}} = \frac{{{\rho ^2}{{\left( {\sin \theta } \right)}^2}}}{{1 + {{\left( {\sin \theta } \right)}^2}}}\frac{{\sin \left( {2\theta } \right)}}{2} \hfill \\ \hfill \\ - \frac{{{\rho ^2}}}{2} \leqslant \frac{{{\rho ^2}{{\left( {\sin \theta } \right)}^2}}}{{1 + {{\left( {\sin \theta } \right)}^2}}}\frac{1}{2} \leqslant f\left( {\rho ,\theta } \right) \leqslant \frac{{{\rho ^2}{{\left( {\sin \theta } \right)}^2}}}{{1 + {{\left( {\sin \theta } \right)}^2}}}\frac{1}{2} \leqslant \frac{{{\rho ^2}}}{2} \hfill \\ \end{gathered} \] since: \[\large - \frac{1}{2} \leqslant \frac{{\sin \left( {2\theta } \right)}}{2} \leqslant \frac{1}{2}\]

OpenStudy (michele_laino):

oops.. I have made a typo, here is the right inequality: \[ \large - \frac{{{\rho ^2}}}{2} \leqslant - \frac{{{\rho ^2}{{\left( {\sin \theta } \right)}^2}}}{{1 + {{\left( {\sin \theta } \right)}^2}}}\frac{1}{2} \leqslant f\left( {\rho ,\theta } \right) \leqslant \frac{{{\rho ^2}{{\left( {\sin \theta } \right)}^2}}}{{1 + {{\left( {\sin \theta } \right)}^2}}}\frac{1}{2} \leqslant \frac{{{\rho ^2}}}{2}\]

OpenStudy (michele_laino):

so, it is suffice to choose \(\delta= \sqrt 2 \; \epsilon\), and the relation of limit, is proven

OpenStudy (kenshin):

sqrt(2epsilon) or sqrt(2)*epsilon?

OpenStudy (kenshin):

shouldn't we make p^2/2 < epsilon?

OpenStudy (michele_laino):

it is: \((\sqrt 2) \epsilon\)

OpenStudy (michele_laino):

yes! We have to solve this equation: \[\huge \frac{{{\delta ^2}}}{2} = \epsilon \]

OpenStudy (michele_laino):

sorry: \[\huge \delta = \sqrt {2\epsilon } \]

OpenStudy (kenshin):

cool, thanks alot!

OpenStudy (michele_laino):

:)

OpenStudy (kenshin):

Hi again, just unsure about the part I msged you about how did -p^2 / 2 <= -p^2sin^2(theta)/(1+sin^2(theta)) appear?

OpenStudy (michele_laino):

step by step, I have used this identities: \[\large \begin{gathered} \cos \theta {\left( {\sin \theta } \right)^3} = {\left( {\sin \theta } \right)^2}\sin \theta \cos \theta = {\left( {\sin \theta } \right)^2}\frac{{2\sin \theta \cos \theta }}{2} = \hfill \\ \hfill \\ = {\left( {\sin \theta } \right)^2}\frac{{\sin \left( {2\theta } \right)}}{2} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

oops.. these* identities

OpenStudy (kenshin):

i understand the sin(2theta)/2 part, but how are the rest ?

OpenStudy (michele_laino):

it is the numerator

OpenStudy (kenshin):

(sin(theta))^2 / (1+sin(theta)^2), how did that become a 1?

OpenStudy (michele_laino):

here are more steps: \[\large \frac{{{\rho ^2}{{\left( {\sin \theta } \right)}^2}}}{{1 + {{\left( {\sin \theta } \right)}^2}}}\frac{{\sin \left( {2\theta } \right)}}{2} \leqslant \frac{{{\rho ^2}{{\left( {\sin \theta } \right)}^2}}}{{{{\left( {\sin \theta } \right)}^2}}}\frac{{\sin \left( {2\theta } \right)}}{2} = {\rho ^2}\frac{{\sin \left( {2\theta } \right)}}{2}\]

OpenStudy (kenshin):

oh gotcha, thanks, so it's not like sin(theta)^2 / (1+sin(theta)^2) can equal 1 or anything, it's just smaller or equal to.

OpenStudy (michele_laino):

yes!